OpenStudy (anonymous):

What is the expression in factored form? 6x^2 + 4x?? help please?

OpenStudy (anonymous):

\[ 6x^2 + 4x = 2 \times 3 \times x \times x+ 2 \times 2 \times x=\] \[= 2 \times x( 3 \times x+ 2)= 2x( 3 x+ 2)\] @lexi503

OpenStudy (anonymous):

thanks sooo much!!:D